A company has unlevered cash flow in years 1-5: $10, $12.5, $13.5, $14.0, $15.0. What is the value of the 5 year projection period assuming a 14.0% discount rate and a 2% perpetuity rate?
Answer #1

We need to calculate the terminal value of the cash flows after year 5.

The terminal value can be given as:

Terminal Value=CF1/(r-g) where CF1 is the next year cash flow
r is the required return and g is the growth rate

Terminal value =15*(1+0.02)/(0.14-0.02)
=$127.50

Hence year 5 cash flows shall have added terminal value also

Year 5 cash flow=$15+$127.50=$142.50

Year Cash flows DF PV
1 $    10.00 0.877193 $       8.77
2 $    12.50 0.769468 $       9.62
3 $    13.50 0.674972 $       9.11
4 $    14.00 0.59208 $       8.29
5 $ 142.50 0.519369 $    74.01
Total $ 109.80

DF is the discount factor =1/(1+r)^n

r=0.14

n is the year

Value =$109.80